2. Determine the normalized eigenfunctions of the given problem y + ly = 0, y(0) = 0, y(1)=0)
Answer #1

g + xy =0 -0 cose Į when A=0 - then equation 0 becomes = =0 2) Y = AXTB / Where A and B care © constants ) y(0) =0 al B=0y c coolu u) + d Sindex) - © (0)=0 => 0 = CX1 +0C70 Hence, the equation ® becomes ya d sinler - diff equation ® woro to sCase to when X=-112 Then equation 0 becomes - y_42y zo The Auxillary equation of o m2_42=0 ) m= Il+ y = E eur +F é un o y cos=o -) 0 = E + F — © =) E = -F. Diffneqn ② wor.do n y = Elleen_Fue un as yll=0 - р - Енеч - Ене-